# Joint Fine-tuning and Conversion of Pretrained Speech and Language Models towards Linear Complexity

### v1

@@ 2.2 Distillation and model conversion
Knowledge distillation. Distilling knowledge from teacher models to a student one by matching the output probabilities is a standard approach to compress ML models, including pretrained language models (LMs) (Hinton et al., 2015; Sanh et al., 2019). Intermediate hidden states of the teacher and student can be further aligned by layerwise distillation (Sun et al., 2019; Aguilar et al., 2020). In addition, Brown et al. (2023) have attempted to distill and compress existing pretrained linear-complexity LMs. In these works, parameters from the teacher are also transferred to the student when possible; an example of the direct parameter transfer approach.
@@ 3 Methods
We aim at building a broadly applicable framework to convert an existing pretrained model into a task-specific linear-complexity one. To achieve this goal, we combine parameter transfer and distillation to best reproduce the performance of the target teacher model, a standard transformer model fine-tuned on the target task from the pretrained source teacher model. For parameter transfer, we replace only the attention layers in the teacher with the more efficient sequence-mixing modules we intend to use, forming the student model as a result; most other layers including embeddings, feed-forward, and layernorm will be preserved. Following this approach, the teacher parameters can be used to initialize the student model as much as possible. This constitutes our baseline unguided approach, in which we directly fine-tune the student model with transferred parameters. Instead, by posing loss on the difference of hidden states, we enforce constraints on the deviation of the student’s hidden states from the teacher. In this way, we directly guide each layer in the student to reproduce the behavior of the corresponding teacher layer. Different possible modes of such guidance are illustrated in Figure 1 and introduced below:
@@
    We set the temperature $\beta=2$ following Hinton et al. (2015). The final loss $\mathcal{L}$ is a combination of the cross entropy $\mathcal{L}_{\text{CE}}$, the KL divergence $\mathcal{L}_{\text{KD}}$ between teacher and student outputs in standard KD, and an MSE loss $\mathcal{L}_{\text{LD}}$ between hidden states for layerwise distillation. The three terms are weighted by a set of hyperparameters $\alpha$. Instead, the unguided mode uses $\mathcal{L}_{\text{CE}}$ only. More specifically, as the target sequence-mixing module (e.g., RNN) is nevertheless distinct from standard attention, it can be more difficult to reproduce the hidden states right after the sequence-mixing layer. Instead, we extract the hidden states after the processing of the following feed-forward layer as ${\bm{H}}$ for layerwise distillation.
@@ 4.1 Model configuration
  • Language modeling by Pythia $\rightarrow$ Mamba: We only replace each attention layer with a Mamba mixer of the same hidden size, initialized under the Mamba scheme.
  • Speech processing by Wav2Vec2 $\rightarrow$ Mamba2: Wav2Vec2 is a bidirectional model, while Mamba2 is unidirectional. Hence we build a bidirectional Mamba2 similar to the one by Zhang et al. (2024b). Each attention layer is replaced by two Mamba2 mixers of the same hidden size, but with expand factor = 1, as shown in Figure 2.
@@
In this way, all the models have roughly the same number of parameters after conversion. Classification tasks are carried out by adding a linear prediction head after mean pooling. The modeling and training configuration under each scenario is further introduced below, and more details including hyperparameters are provided in Appendix A.
@@ 4.2 Language processing
The results are given in Table 1. Compared with our fine-tuned standard RoBERTa, the re-pretrained Linformer has a slight drop in performance (1.3% on average). This is typical and can be explained by the reduced expressivity compared to the standard attention. While without extensive pretraining, unguided conversion with only parameter transfer leads to much lower accuracy and severe training instability. Despite much lower computational costs, such performance regression is unacceptable. Fortunately, the regression can be alleviated using CALD. By simply using the target guided approach, the performance drop can be reduced to merely 0.7% on average compared to pretrained Linformer. With the trajectory guided approach, the CALD model can even reach better results (+0.2% on average) but without any pretraining. The approximated waypoint guided approach leads to only slightly lower results, but still better than the target guided one. The hybrid approach leads to lower results compared to target guided, indicating that the guidance is helpful even after early training updates, which matches the rather low unguided results. In addition, we try to use the target guided approach but initialized from the teacher source parameters, which shows slightly lower results. All these results indicate that the behavior transfer by CALD is effective and critical for the joint conversion and fine-tuning in this scenario to minimize the performance regression. Particularly, the trajectory or waypoint guided approaches are helpful.
@@ 4.4 Speech processing
However, the waypoint guided approach is found to be not helpful or even harmful; we choose not to proceed with the more precise trajectory guided approach. Our assumption is that the trajectory guided approach helps retain the knowledge in the hidden states of the pretrained model. Hence We hypothesize that the hidden states of the Wav2Vec2 model undergo significant shift during fine-tuning. As a result, the hidden states of the original pretrained model or those in the early phase of fine-tuning are less useful for the target task. To examine the hypothesis, for each model, we compute the average cosine distance of the hidden states produced by the source teacher and the waypoints, using 100 random data samples. As shown in Figure 3, as for fine-tuning RoBERTa on NLP tasks, the cosine distance is relatively small and increases gradually. While the distance is much larger during the fine-tuning of Wav2Vec2 on speech tasks, even at the early phase of fine-tuning. In fact, the distance has already reached 0.372 after only 2,000 steps of fine-tuning on TED-LIUM. Therefore, our hypothesis is supported that the trajectory guided approach works better in the cases when there is limited hidden state shift during fine-tuning, and when the pretrained model features remain relevant for the target task. While the condition is not satisfied for speech tasks.
@@ Appendix A Implementation details
Std. RoBERTa LR 1e-5 3e-5 1e-5 1e-5
Unguided LR 1e-4 1e-4 1e-4 3e-4
@@
We build the models as mentioned above, and train the models using the AdamW optimizer. On Mamba/Mamba2 models, we also find it important to apply layer normalization on hidden states prior to computing the layerwise distillation loss, i.e. we take the hidden states after the layernorm operation at the following layer. Otherwise the hidden states will have a rather wide range and bring training instability. We only use output distillation by setting $\alpha_{KD}=1$ in ASR experiments. On other experiments, we find that the distillation on output probabilities actually leads to suboptimal results, and thus set $\alpha_{KD}=0$. We always use $\alpha_{CE}=1$. As for the hybrid mode of distillation, we start with the optimal target-guided configuration, but switch to unguided mode when the loss is close to converge, generally at around 30% of the total training steps.

### v2

@@ 2.2 Distillation and model conversion
Knowledge distillation. Distilling knowledge from teacher models to a student one by matching the output probabilities is a standard approach to compress ML models, including pretrained language models (LMs) (Hinton et al., 2015; Sanh et al., 2019). Intermediate hidden states of the teacher and student can be further aligned by layerwise distillation (Sun et al., 2019; Aguilar et al., 2020). In addition, Brown et al. (2023) have attempted to distill and compress existing pretrained linear-complexity LMs. In these works, parameters from the teacher are also transferred to the student when possible, hence also represent the direct parameter transfer approach.
@@ 3 Methods
We aim at building a broadly applicable framework to convert an existing pretrained model into a task-specific linear-complexity one. To achieve this goal, we combine parameter transfer and distillation to best reproduce the performance of the target teacher model, a standard transformer model fine-tuned on the target task from the pretrained source teacher model. For parameter transfer, we replace only the attention layers in the teacher with the more efficient sequence-mixing modules we intend to use, forming the student model as a result; most other layers including embeddings, feed-forward, and layernorm will be preserved. Following this approach, the teacher parameters can be used to initialize the student model as much as possible. This constitutes our baseline unguided approach, in which we directly fine-tune the student model with transferred parameters. Instead, by posing loss on the difference of hidden states, we enforce constraints on the deviation of the student’s hidden states from the teacher. In this way, we directly guide each layer in the student to reproduce the behavior of the corresponding teacher layer. In addition, we consider different ways or modes of such guidance, introduced below. These modes are also illustrated in Figure 1, and pseudo-codes for the different modes of training algorithms are available in Appendix B.
@@
    We set the temperature $\beta=2$ following Hinton et al. (2015). The final loss $\mathcal{L}$ is a combination of the cross entropy $\mathcal{L}_{\text{CE}}$, the KL divergence $\mathcal{L}_{\text{KD}}$ between teacher and student outputs in standard KD, and an MSE loss $\mathcal{L}_{\text{LD}}$ between hidden states for layerwise distillation. The three terms are weighted by a set of hyperparameters $\alpha$. Instead, the unguided mode uses $\mathcal{L}_{\text{CE}}$ only. More specifically, as the target sequence-mixing module (e.g., RNN) is nevertheless distinct from standard attention, it can be more difficult to reproduce the hidden states right after the sequence-mixing layer. Instead, we extract the hidden states after the processing of the following feed-forward layer as ${\bm{H}}$ for layerwise distillation. As for tasks other than classification, we will replace the $\mathcal{L}_{\text{CE}}$ term with the corresponding loss function, e.g. CTC loss for ASR tasks.
@@ 4.1 Model configuration
  • Language modeling by Pythia $\rightarrow$ Mamba: We only replace each attention layer with a Mamba mixer of the same hidden size. The mixer is initialized using the specific Mamba state-space scheme to best preserve the past memory (Gu & Dao, 2023).
  • Speech processing by Wav2Vec2 $\rightarrow$ Mamba2: Wav2Vec2 is a bidirectional model, while Mamba2 is unidirectional. Hence we build a bidirectional Mamba2 similar to the one by Zhang et al. (2024b). Each attention layer is replaced by two Mamba2 mixers of the same hidden size, but with expand factor = 1, as shown in Figure 2. The mixer is similarly initialized with the specific Mamba2 state-space scheme (Dao & Gu, 2024).
@@
In this way, all the models have roughly the same number of parameters after conversion. Classification tasks are carried out by adding a linear prediction head after mean pooling. The modeling and training configuration under each scenario is further introduced below, and more details including hyperparameters are provided in Appendix A. Additional details on the computational costs are given in Appendix C.
@@ 4.2 Language processing
The results are given in Table 1. Compared with our fine-tuned standard RoBERTa, the re-pretrained Linformer has a slight drop in performance (1.3% on average). This is typical and can be explained by the reduced expressivity compared to the standard attention. While without extensive pretraining, unguided conversion with only parameter transfer leads to much lower accuracy and severe training instability. Despite much lower computational costs, such performance regression is unacceptable. Fortunately, the regression can be alleviated using CALD. By simply using the target guided approach, the performance drop can be reduced to merely 0.7% on average compared to pretrained Linformer. With the trajectory guided approach, the CALD model can even reach better results (+0.2% on average) but without any pretraining. The approximated waypoint guided approach leads to only slightly lower results, but still better than the target guided one. The hybrid approach leads to lower results compared to target guided, indicating that the guidance is helpful even after early training updates, which matches the rather low unguided results; this is further discussed below in Section 4.4. In addition, we try to use the target guided approach but initialized from the teacher source parameters. Since the distillation should be easier if the student model is initialized with parameters exactly from the teacher model (i.e. the teacher target), these Src. init. models show lower results as we expected. All these results indicate that the behavior transfer by CALD is effective and critical for the joint conversion and fine-tuning in this scenario to minimize the performance regression. Particularly, the trajectory or waypoint guided approaches are helpful.
@@ 4.4 Speech processing
However, the waypoint guided approach is found to be not helpful or even harmful; we choose not to proceed with the more precise trajectory guided approach. Our assumption is that the trajectory guided approach helps retain the knowledge in the hidden states of the pretrained model. Hence We hypothesize that the hidden states of the Wav2Vec2 model undergo significant shift during fine-tuning. As a result, the hidden states of the original pretrained model or those in the early phase of fine-tuning are less useful for the target task. To examine the hypothesis, for each model, we compute the average cosine distance of the hidden states produced by the source teacher and the waypoints, using 100 random data samples. As shown in Figure 3, as for fine-tuning RoBERTa on NLP tasks, the cosine distance is relatively small and increases gradually. While the distance is much larger during the fine-tuning of Wav2Vec2 on speech tasks, even at the early phase of fine-tuning. In fact, the distance has already reached 0.372 after only 2,000 steps of fine-tuning on TED-LIUM. Therefore, our hypothesis is supported that the trajectory guided approach works better in the cases when there is limited hidden state shift during fine-tuning, and when the pretrained model features remain relevant for the target task. While the condition is not satisfied for speech tasks, in which case allowing the model to deviate more from the initial behavior using the hybrid approach will be beneficial.
In addition, we have also attempted to train the models that are unguided and without any parameter transfer. Such a large Mamba2-based speech model trained on the target task from scratch is rather unstable and fails to converge well. This highlights the necessity of pretraining in such large models.
@@ Appendix A Implementation details
We build the models as mentioned above, and train the models using the AdamW optimizer. On Mamba/Mamba2 models, we also find it important to apply layer normalization on hidden states prior to computing the layerwise distillation loss, i.e. we take the hidden states after the layernorm operation at the following layer. Otherwise the hidden states will have a rather wide range and bring training instability. We only use output distillation by setting $\alpha_{KD}=1$ in ASR experiments. On other experiments, we find that the distillation on output probabilities actually leads to suboptimal results, and thus set $\alpha_{KD}=0$. We always use $\alpha_{CE}=1$. As for the hybrid mode of distillation, we start with the optimal target guided configuration, but switch to unguided mode when the loss is close to converge, generally at around 30% of the total training steps.
@@
Std. RoBERTa LR 3e-5 3e-5 3e-5 1e-5
Unguided LR 1e-5 1e-5 3e-5 1e-5
Share mode KV Layer KV KV
@@
Appendix B Distillation algorithms
Here we provide the pseudo-code to describe the training algorithms under different modes of distillation to facilitate understanding. Please refer to Section 3 for the notation scheme we use, including the model output terms and loss functions.
Algorithm 1 CALD training under the target guided or hybrid mode.
0:  mode: training mode (hybrid or target guided); $T_{h}$: number of guided steps in the hybrid mode; $T$: total number of training steps; dataloader: loader of training data; $\alpha_{CE}$, $\alpha_{KD}$, $\alpha_{LD}$: hyperparameters for loss term scaling.
1:  Build and initialize the teacher model $M_{T}$ and the student model $M_{S}$ with the teacher target model.
2:  Replace the attention layers in $M_{S}$.
3:  Build the optimizer for $M_{S}$.
4:  for $i=1$ to $T$ do
[…]
6:     $(y^{(s)},H^{(s)})\leftarrow M_{S}(x)$
7:     if mode = Hybrid and $i>T_{h}$ then
8:        $\mathcal{L}\leftarrow\alpha_{\text{CE}}\cdot\mathcal{L}_{\text{CE}}(y^{(s)},y)$
9:     else
10:        $(y^{(t)},H^{(t)})\leftarrow M_{T}(x)$
11:        $\mathcal{L}\leftarrow\alpha_{\text{CE}}\cdot\mathcal{L}_{\text{CE}}(y^{(s)},y)+\alpha_{\text{KD}}\cdot\mathcal{L}_{\text{KD}}(y^{(s)},y^{(t)})+\alpha_{\text{LD}}\cdot\mathcal{L}_{\text{LD}}(H^{(s)},H^{(t)})$
12:     end if
13:     $\mathcal{L}$.backward()
14:     Update $M_{S}$ parameters using the optimizer.
15:  end for
Algorithm 2 CALD training under the trajectory guided mode.
0:  $T_{u}$: interval between teacher model updates; $T$: total number of training steps; dataloader: loader of training data; $\alpha_{CE}$, $\alpha_{KD}$, $\alpha_{LD}$: hyperparameters for loss term scaling.
1:  Build and initialize the teacher model $M_{T}$ and the student model $M_{S}$ with the teacher source model (the original pretrained model).
2:  Replace the attention layers in $M_{S}$.
3:  Build the optimizer for $M_{S}$ and $M_{T}$.
4:  for $i=1$ to $T$ do
[…]
6:     $(y^{(t)},H^{(t)})\leftarrow M_{T}(x)$
7:     if $i$ mod $T_{u}$ = 0 then
8:        $\mathcal{L}_{T}=\mathcal{L}_{\text{CE}}(y^{(t)},y)$.
9:        $\mathcal{L}_{T}$.backward()
10:        Update $M_{T}$ parameters using the optimizer.
11:     end if
12:     $(y^{(s)},H^{(s)})\leftarrow M_{S}(x)$
13:     $\mathcal{L}\leftarrow\alpha_{\text{CE}}\cdot\mathcal{L}_{\text{CE}}(y^{(s)},y)+\alpha_{\text{KD}}\cdot\mathcal{L}_{\text{KD}}(y^{(s)},y^{(t)})+\alpha_{\text{LD}}\cdot\mathcal{L}_{\text{LD}}(H^{(s)},H^{(t)})$
14:     $\mathcal{L}$.backward()
15:     Update $M_{S}$ parameters using the optimizer.
16:  end for
Algorithm 3 CALD training under the waypoint guided mode.
0:  $W$: list of waypoint models; $T_{w}$: interval between waypoint switching; $T$: total number of training steps; dataloader: loader of training data; $\alpha_{CE}$, $\alpha_{KD}$, $\alpha_{LD}$: hyperparameters for loss term scaling.
1:  Build and initialize the teacher model $M_{T}$ with $W_{1}$, $w_{i}\leftarrow 1$.
2:  Build and initialize the student model $M_{S}$ from the teacher source model (the original pretrained model).
3:  Replace the attention layers in $M_{S}$.
4:  Build the optimizer for $M_{S}$.
5:  for $i=1$ to $T$ do
[…]
7:     $(y^{(s)},H^{(s)})\leftarrow M_{S}(x)$
8:     $(y^{(t)},H^{(t)})\leftarrow M_{T}(x)$
9:     $\mathcal{L}\leftarrow\alpha_{\text{CE}}\cdot\mathcal{L}_{\text{CE}}(y^{(s)},y)+\alpha_{\text{KD}}\cdot\mathcal{L}_{\text{KD}}(y^{(s)},y^{(t)})+\alpha_{\text{LD}}\cdot\mathcal{L}_{\text{LD}}(H^{(s)},H^{(t)})$
10:     $\mathcal{L}$.backward()
11:     Update $M_{S}$ parameters using the optimizer.
12:     if $i$ mod $T_{w}$ = 0 then
13:        $w_{i}\leftarrow w_{i}+1$
14:        Re-initialize $M_{T}$ with $W_{w_{i}}$.
15:     end if
16:  end for
Appendix C Computational costs
The proposed method is presented as a more efficient alternative to re-doing the whole pretraining process on the target architecture, as the model pretraining is notoriously expensive and generally infeasible with academic computation. For example, Wav2Vec2-large pretraining took 5.2 days on 128 V100 GPUs (Baevski et al., 2020), while our target guided CALD experiment to convert Wav2Vec2-large into Mamba2 for ASR takes only 1.6 days on a single RTX3090 with just 24GB memory. Without the computation on the teacher model, the unguided training will be roughly 40% faster, but the accuracy degradation is unacceptable. The hybrid approach is faster owing to more than half of the training steps being under the unguided mode, while reaching better performance than the target guided one. The waypoint guided model doesn’t involve any extra computation compared to the target guided one, and hence takes roughly the same time. In NLP experiments, the trajectory guided models enjoy better performance at the cost of more computation. For example, on QNLI with layer-shared Linformer projection, the target or waypoint guided training takes around 3.4 hours under our training configuration with a single RTX3090, while the trajectory guided training takes 3.5 hours ($T_{u}=5$) to 4.0 hours ($T_{u}=1$). The unguided training takes around 3.0 hours. To sum up, the CALD approach saves much computation compared to re-pretraining, and the extra cost compared to the unguided training is limited.
We also provide a brief evaluation on the inference speed of transformer-based versus Mamba2-based ASR models we used in Section 4.4, as shown in Figure 5. With asymptotic linear time complexity, the Mamba2 model will be much faster given sufficiently long inputs compared to the quadratic time transformers, although the actual speedup depends on the exact implementation; an investigation on this topic falls out of the scope of this paper.
Figure 5: Time costs for Mamba2 and transformer-based models performing ASR on audio samples with different lengths, averaged by 5 runs on a single RTX3090.